# Linear Mode Connectivity in Differentiable Tree Ensembles

### v1

@@ Abstract
Linear Mode Connectivity (LMC) refers to the phenomenon that performance remains consistent for linearly interpolated models in the parameter space. For independently optimized model pairs from different random initializations, achieving LMC is considered crucial for validating the stable success of the non-convex optimization in modern machine learning models and for facilitating practical parameter-based operations such as model merging. While LMC has been achieved for neural networks by considering the permutation invariance of neurons in each hidden layer, its attainment for other models remains an open question. In this paper, we first achieve LMC for soft tree ensembles, which are tree-based differentiable models extensively used in practice. We show the necessity of incorporating two invariances: subtree flip invariance and splitting order invariance, which do not exist in neural networks but are inherent to tree architectures, in addition to permutation invariance of trees. Moreover, we demonstrate that it is even possible to exclude such additional invariances while keeping LMC by designing decision list-based tree architectures, where such invariances do not exist by definition. Our findings indicate the significance of accounting for architecture-specific invariances in achieving LMC.
@@ 1 Introduction
A non-trivial empirical characteristic of modern machine learning models trained using gradient methods is that models trained from different random initializations could become functionally almost equivalent, even though their parameter representations differ. If the outcomes of all training sessions converge to the same local minima, this empirical phenomenon can be understood. However, considering the complex non-convex nature of the loss surface, the optimization results are unlikely to converge to the same local minima. In recent years, particularly within the context of neural networks, the transformation of model parameters while preserving functional equivalence has been explored by considering the permutation invariance of neurons in each hidden layer [1, 2]. Notably, only a slight performance degradation has been observed when using weights derived through linear interpolation between permuted parameters obtained from different training processes [3, 4]. This demonstrates that the trained models reside in different, yet functionally equivalent, local minima. This situation is referred to as Linear Mode Connectivity (LMC) [5]. From a theoretical perspective, LMC is crucial for supporting the stable and successful application of non-convex optimization. In addition, LMC also holds significant practical importance, enabling techniques such as model merging [6, 7] by weight-space parameter averaging.
@@
Although neural networks are most extensively studied among the models trained using gradient methods, other models also thrive in real-world applications. A representative is tree ensemble models, such as random forests [8]. While they are originally trained by not gradient but greedy algorithms, differentiable soft tree ensembles, which learn parameters of the entire model through gradient-based optimization, have recently been actively studied. Not only empirical studies regarding accuracy and interpretability [9, 10, 11], but also theoretical analyses have been performed [12, 13]. Moreover, the differentiability of soft trees allows for integration with various deep learning methodologies, including fine-tuning [14], dropout [15], and various stochastic gradient descent methods [16, 17]. Furthermore, the soft tree represents the most elementary form of a hierarchical mixture of experts [18, 19, 20]. Investigating soft tree models not only advances our understanding of this particular structure but also contributes to broader research into essential technological components critical for the development of large-scale language models [21].
@@
Figure 1: A representative experimental result on the MiniBooNE [22] dataset (left) and conceptual diagram of the LMC for tree ensembles (right).
@@
A research question that we tackle in this paper is: “Can LMC be achieved for soft tree ensembles?”. Our empirical results, which are highlighted with a green line in the top left panel of Figure 1, clearly show that the answer is “Yes”. This plot shows the variation in test accuracy when interpolating weights of soft oblivious trees, perfect binary soft trees with shared parameters at each depth, trained from different random initializations. The green line is obtained by our method introduced in this paper, where there is almost zero performance degradation. Furthermore, as shown in the bottom left panel of Figure 1, the performance can even improve when interpolating between models trained on split datasets.
@@
Moreover, we demonstrate that it is possible to exclude such additional invariances while preserving LMC by modifying tree architectures. We realize such an architecture based on a decision list, a binary tree structure where branches extend in only one direction. By designating one of the terminal leaves as an empty node, we introduce a customized decision list that omits both subtree flip invariance and splitting order invariance, and empirically show that this can achieve LMC by considering only tree permutation invariance. Since incorporating additional invariances is computationally expensive, we can efficiently perform weight-space averaging in model merging on our customized decision lists.
@@
  • First achievement of LMC for tree ensembles with accounting for additional invariances beyond tree permutation.
  • Development of a decision list-based tree architecture that does not involve the additional invariances.
  • A thorough empirical investigation of LMC across various tree architectures, invariances, and real-world datasets.
@@ 2.1 Linear Mode Connectivity
Let us consider two models, $A$ and $B$, that have the same architecture. In the context of evaluating LMC, the concept of a “barrier” is frequently used [4, 23]. Let $\boldsymbol{\Theta}_{A},\boldsymbol{\Theta}_{B}\in\mathbb{R}^{P}$ be vectorized parameters of models $A$ and $B$, respectively, for $P$ parameters. Assume that $\mathcal{C}:\mathbb{R}^{P}\rightarrow\mathbb{R}$ measures the performance of the model, such as accuracy, given its parameter vector. If higher values of $\mathcal{C}(\cdot)$ mean better performance, the barrier between two parameter vectors $\boldsymbol{\Theta}_{A}$ and $\boldsymbol{\Theta}_{B}$ is defined as:
@@
Several techniques have been developed to reduce barriers by transforming parameters while preserving functional equivalence. Two main approaches are activation matching (AM) and weight matching (WM). AM takes the behavior of model inference into account, while WM simply compares two models using their parameters. The validity of both AM and WM has been theoretically supported [24]. Numerous algorithms are available for implementing AM and WM. For instance, [4] uses a formulation based on the Linear Assignment Problem (LAP) to find suitable permutations, while [23] employs a differentiable formulation that allows for the optimization of permutations using gradient-based methods.
Existing research has focused exclusively on neural network architectures such as multi-layer perceptrons (MLP) and convolutional neural networks (CNN). No study has been conducted from the perspective of linear mode connectivity for soft tree ensembles.
@@ 2.2 Soft Tree Ensemble
Unlike typical hard decision trees, which explicitly determine the data flow to the right or left at each splitting node, soft trees represent the proportion of data flowing to the right or left as continuous values between 0 and 1. This approach enables a differentiable formulation.
We use a $\operatorname{sigmoid}$ function, $\sigma:\mathbb{R}\to(0,1)$ to formulate a function $\mu_{m,\ell}(\boldsymbol{x}_{i},\boldsymbol{w}_{m},\boldsymbol{b}_{m}):\mathbb{R}^{F}\times\mathbb{R}^{F\times\mathcal{N}}\times\mathbb{R}^{1\times\mathcal{N}}\to(0,1)$ that represents the proportion of the $i$th data point $\boldsymbol{x}_{i}$ flowing to the $\ell$th leaf of the $m$th tree as a result of soft splittings:
@@
To classify $C$ categories, the output of the $m$th tree is computed by the function $f_{m}:\mathbb{R}^{F}\times\mathbb{R}^{F\times\mathcal{N}}\times\mathbb{R}^{1\times\mathcal{N}}\times\mathbb{R}^{C\times\mathcal{L}}\to\mathbb{R}^{C}$ as sum of the leaf parameters $\boldsymbol{\pi}_{m,\ell}$ weighted by the outputs of $\mu_{m,\ell}(\boldsymbol{x}_{i},\boldsymbol{w}_{m},\boldsymbol{b}_{m})$:
$f_{m}(\boldsymbol{x}_{i},\boldsymbol{w}_{m},\boldsymbol{b}_{m},\boldsymbol{\pi}_{m})=\sum_{\ell=1}^{\mathcal{L}}\boldsymbol{\pi}_{m,\ell}\mu_{m,\ell}(\boldsymbol{x}_{i},\boldsymbol{w}_{m},\boldsymbol{b}_{m}),$ (3)
where $\mathcal{L}$ is the number of leaves in a tree. By combining this function for $M$ trees, we realize the function $f:\mathbb{R}^{F}\times\mathbb{R}^{M\times F\times\mathcal{N}}\times\mathbb{R}^{M\times 1\times\mathcal{N}}\times\mathbb{R}^{M\times C\times\mathcal{L}}\to\mathbb{R}^{C}$ as an ensemble model consisting of $M$ trees:
@@
with the parameters $\boldsymbol{w}=(\boldsymbol{w}_{1},\dots,\boldsymbol{w}_{M})$, $\boldsymbol{b}=(\boldsymbol{b}_{1},\dots,\boldsymbol{b}_{M})$, and $\boldsymbol{\pi}=(\boldsymbol{\pi}_{1},\dots,\boldsymbol{\pi}_{M})$ being randomly initialized.
Despite the apparent differences, there are correspondences between MLPs and soft tree ensemble models. The formulation of a soft tree ensemble with $D=1$ is:
$f(\boldsymbol{x}_{i},\boldsymbol{w},\boldsymbol{b},\boldsymbol{\pi})$ $=\sum_{m=1}^{M}\Bigl{(}\,\sigma(\boldsymbol{w}_{m,1}^{\top}\boldsymbol{x}_{i}+b_{m,1})\boldsymbol{\pi}_{m,1}+(1-\sigma(\boldsymbol{w}_{m,1}^{\top}\boldsymbol{x}_{i}+b_{m,1}))\boldsymbol{\pi}_{m,2}\,\Bigr{)}$
$=\sum_{m=1}^{M}\Bigl{(}\,(\boldsymbol{\pi}_{m,1}-\boldsymbol{\pi}_{m,2})\sigma(\boldsymbol{w}_{m,1}^{\top}\boldsymbol{x}_{i}+b_{m,1})+\boldsymbol{\pi}_{m,2}\,\Bigr{)}.$ (5)
When we consider the correspondence between $\boldsymbol{\pi}_{m,1}-\boldsymbol{\pi}_{m,2}$ in tree ensembles and second layer weights in the two-layer perceptron, the tree ensembles model matches to the two-layer perceptron. It is clear from the formulation that the permutation of hidden neurons in a neural network corresponds to the rearrangement of trees in a tree ensemble.
@@ 3 Invariances Inherent to Tree Ensembles
3.1 Parameter modification processes that maintains functional equivalence in tree ensembles
First, we clarify what invariances should be considered for tree ensembles, which are expected to reduce the barrier significantly if taken into account. When we consider perfect binary trees, there are three types of invariance:
Figure 2: (a) Subtree flip invariance. (b) Splitting order invariance for an oblivious tree.
  • Tree permutation invariance. In Equation (4), the behavior of the function does not change even if the order of the $M$ trees is altered. This corresponds to the permutation of internal nodes in neural networks, which has been a subject of active interest in previous studies on LMC.
  • Subtree flip invariance. When the left and right subtrees are swapped simultaneously with the inversion of the inequality sign at the split, the functional behavior remains unchanged, which we refer to subtree flip invariance. Figure 2(a) presents a schematic diagram of this invariance, which is not found in neural networks but is unique to binary tree-based models. Since $\sigma(-c)=1-\sigma(c)$ for $c\in\mathbb{R}$ due to the symmetry of $\operatorname{sigmoid}$, the inversion of the inequality is achieved by inverting the signs of $\boldsymbol{w}_{m,n}$ and $b_{m,n}$. [25] also focused on the sign of weights, but in a different way from ours. They pay attention to the amount of change from the parameters at the start of fine-tuning, rather than discussing the sign of the parameters.
  • Splitting order invariance. Oblivious trees share parameters at the same depth, which means that the decision boundaries are straight lines without any bends. With this characteristic, even if the splitting rules at different depths are swapped, functional equivalence can be achieved if the positions of leaves are also swapped appropriately as shown in Figure 2(b). This invariance does not exist for non-oblivious perfect binary trees without parameter sharing, as the behavior of the decision boundary varies depending on the splitting order.
Note that MLPs also have an additional invariance beyond just permutation. Particularly in MLPs that employ ReLU as an activation function, the output of each layer changes linearly with a zero crossover. Therefore, it is possible to modify parameters without changing functional behavior by multiplying the weights in one layer by a constant and dividing the weights in the previous layer by the same constant. However, since the soft tree is based on the $\operatorname{sigmoid}$ function, this invariance does not apply. Previous studies [3, 4, 23] have consistently achieved significant reductions in barriers without accounting for this scale invariance. One potential reason is that changes in parameter scale are unlikely due to the nature of optimization via gradient descent. Conversely, when we consider additional invariances inherent to trees, the scale is equivalent to the original parameters.

### v2

@@ Abstract
Linear Mode Connectivity (LMC) refers to the phenomenon that performance remains consistent for linearly interpolated models in the parameter space. For independently optimized model pairs from different random initializations, achieving LMC is considered crucial for understanding the stable success of the non-convex optimization in modern machine learning models and for facilitating practical parameter-based operations such as model merging. While LMC has been achieved for neural networks by considering the permutation invariance of neurons in each hidden layer, its attainment for other models remains an open question. In this paper, we first achieve LMC for soft tree ensembles, which are tree-based differentiable models extensively used in practice. We show the necessity of incorporating two invariances: subtree flip invariance and splitting order invariance, which do not exist in neural networks but are inherent to tree architectures, in addition to permutation invariance of trees. Moreover, we demonstrate that it is even possible to exclude such additional invariances while keeping LMC by designing decision list-based tree architectures, where such invariances do not exist by definition. Our findings indicate the significance of accounting for architecture-specific invariances in achieving LMC.
@@ 1 Introduction
A non-trivial empirical characteristic of modern machine learning models trained using gradient methods is that models trained from different random initializations could achieve nearly identical performance, even though their parameter representations differ. This empirical phenomenon can be understood if the outcomes of all training sessions converge to the same local minima. However, considering the complex non-convex nature of the loss surface, the optimization results are unlikely to converge to the same local minima. In recent years, particularly within the context of neural networks, the transformation of model parameters while preserving functional equivalence has been explored by considering the permutation invariance of neurons in each hidden layer [1, 2]. Notably, only a slight performance degradation has been observed when using weights derived through linear interpolation between permuted parameters obtained from different training processes [3, 4]. This demonstrates that the trained models reside in different, yet equivalent, local minima. This situation is referred to as Linear Mode Connectivity (LMC) [5]. From a theoretical perspective, LMC is crucial for understanding the stable and successful application of non-convex optimization. As noted by [3] and [4], achievement of LMC suggests that loss landscapes often contain (nearly) a single basin after accounting for all possible invariances, which can be an intuitive reason for the robustness of gradient methods to different random initialization and data batch orders. In addition, LMC also holds significant practical importance, enabling techniques such as model merging [6, 7].
@@
Although neural networks have been studied most extensively studied among the models trained using gradient methods, other models also thrive in real-world applications. A representative is decision tree ensemble models, such as random forests [8]. A decision tree ensemble makes predictions by combining the outputs of multiple trees that recursively split the data into subsets at each node and make final predictions at their leaves. While they are originally trained by greedy algorithms, not by gradient algorithms, their differentiable variant — called soft tree ensembles, which learn parameters of the entire model through gradient-based optimization — have recently been actively studied. Not only empirical studies regarding accuracy and interpretability [9, 10, 11], but also theoretical analyses have been performed [12, 13]. Moreover, the differentiability of soft trees allows for integration with various deep learning methodologies, including fine-tuning [14], dropout [15], and various stochastic gradient descent methods [16, 17]. Furthermore, the soft tree represents the most elementary form of a hierarchical mixture of experts [18]. Investigating soft tree models not only advances our understanding of this particular structure but also contributes to broader research on the key components that are essential to developing large-scale models [19].
@@
Figure 1: A representative experimental result on the MiniBooNE [20] dataset (left) and conceptual diagram of LMC for tree ensembles (right).
A research question that we tackle in this paper is: “Can LMC be achieved for soft tree ensembles?”. While achieving LMC has advanced the understanding of non-convex optimization and the use of model merging in neural networks, it has yet to be explored in tree ensemble models. The reasons behind achieving LMC, even in neural networks, are not fully understood, and it is also unclear whether LMC can be realized in soft tree ensembles, given their distinct architectures. Thus, our contribution of examining LMC in soft tree ensembles provides not only novel insights and techniques for tree ensemble models but also broadens the understanding of the LMC phenomenon by introducing perspectives beyond neural networks for the first time.
@@
Our results, which are highlighted with a green line in the top left panel of Figure 1, clearly show that the answer to our research question is “Yes”. This plot shows the variation in test accuracy when interpolating weights of soft oblivious trees, perfect binary soft trees with shared parameters at each depth, trained from different random initializations. The green line is obtained by the method introduced in this paper, where there is almost zero performance degradation. Furthermore, as shown in the bottom left panel of Figure 1, the performance can even improve when interpolating between models trained on split datasets.
@@
Moreover, we demonstrate that it is possible to exclude such additional invariances while preserving LMC by modifying tree architectures. We realize such an architecture based on a decision list, a binary tree structure where branches extend in only one direction. By designating one of the terminal leaves as an empty node, we introduce a customized decision list that omits both subtree flip invariance and splitting order invariance, and empirically show that this can achieve LMC by considering only tree permutation invariance. Since incorporating additional invariances is computationally expensive, we can efficiently perform model merging on our customized decision lists.
@@
  • First achievement of LMC for tree ensembles with additional invariances beyond tree permutation.
  • Development of a decision list-based architecture that does not involve additional invariances.
  • A thorough empirical investigation of LMC across various tree architectures and real-world datasets.
@@ 2.1 Linear Mode Connectivity
Let us consider two models, $A$ and $B$, which have the same architecture. In the context of evaluating LMC, the concept of a “barrier” is frequently used [4, 21]. Let $\boldsymbol{\Theta}_{A}$ and $\boldsymbol{\Theta}_{B}$ be the parameters of models $A$ and $B$, respectively. Their shapes can be defined arbitrarily. In this paper, for tree ensemble models, $\boldsymbol{\Theta}_{A},\boldsymbol{\Theta}_{B}\in\mathbb{R}^{M\times P}$ for the number $P$ of parameters per tree and the number $M$ of trees. Assume that $\mathcal{C}:\mathbb{R}^{M\times P}\rightarrow\mathbb{R}$ measures the performance of the model, such as accuracy. If higher values of $\mathcal{C}(\cdot)$ mean better performance, the barrier between two parameter vectors $\boldsymbol{\Theta}_{A}$ and $\boldsymbol{\Theta}_{B}$ is defined as:
@@
Several techniques have been developed to reduce barriers by transforming parameters while preserving functional equivalence. Two main approaches are activation matching (AM) and weight matching (WM). AM takes the behavior of model inference into account, while WM simply compares two models using their parameters. The validity of both AM and WM has been theoretically supported by [22]. Numerous algorithms are available for implementing AM and WM. For instance, [4] used a formulation based on the Linear Assignment Problem (LAP), also known as finding the minimum-cost matching in bipartite graphs, to determine suitable permutations. [21] employed a differentiable formulation that allows for the optimization of permutations using gradient-based methods.
Existing research has focused exclusively on neural networks such as multi-layer perceptrons (MLP) and convolutional neural networks (CNN). No studies have been conducted for soft tree ensembles.
@@ 2.2 Soft Tree Ensemble
Unlike typical hard decision trees, which explicitly determine the data flow to the right or left at each splitting node, soft trees represent the proportion of data flowing to the right or left as continuous values between 0 and 1. This approach enables a differentiable formulation. We use a $\operatorname{sigmoid}$ function, $\sigma:\mathbb{R}\to(0,1)$ to formulate a function $\mu_{m,\ell}(\boldsymbol{x}_{i},\boldsymbol{w}_{m},\boldsymbol{b}_{m}):\mathbb{R}^{F}\times\mathbb{R}^{F\times\mathcal{N}}\times\mathbb{R}^{1\times\mathcal{N}}\to(0,1)$ that represents the proportion of the $i$th data point $\boldsymbol{x}_{i}$ flowing to the $\ell$th leaf of the $m$th tree as a result of soft splittings:
@@
Figure 2: A soft decision tree with a single inner node and two leaf nodes.
To classify $C$ categories, the output of the $m$th tree is computed by the function $f_{m}:\mathbb{R}^{F}\times\mathbb{R}^{F\times\mathcal{N}}\times\mathbb{R}^{1\times\mathcal{N}}\times\mathbb{R}^{C\times\mathcal{L}}\to\mathbb{R}^{C}$ as sum of the leaves $\boldsymbol{\pi}_{m,\ell}$ weighted by the outputs of $\mu_{m,\ell}(\boldsymbol{x}_{i},\boldsymbol{w}_{m},\boldsymbol{b}_{m})$:
$f_{m}(\boldsymbol{x}_{i},\boldsymbol{w}_{m},\boldsymbol{b}_{m},\boldsymbol{\pi}_{m})=\sum_{\ell=1}^{\mathcal{L}}\mu_{m,\ell}(\boldsymbol{x}_{i},\boldsymbol{w}_{m},\boldsymbol{b}_{m})\boldsymbol{\pi}_{m,\ell},$ (3)
where $\mathcal{L}$ is the number of leaves in a tree. To facilitate understanding, the formulation for tree depth is $D=1$ is illustrated in Figure 2.
If $\mu_{m,\ell}(\boldsymbol{x}_{i},\boldsymbol{w}_{m},\boldsymbol{b}_{m})$ takes the value of $1.0$ for one leaf and $0.0$ for the others, the leaf value itself becomes the prediction output, making the model behavior equivalent to that of a standard oblique decision tree [23].
By combining this function for $M$ trees, we realize the function $f:\mathbb{R}^{F}\times\mathbb{R}^{M\times F\times\mathcal{N}}\times\mathbb{R}^{M\times 1\times\mathcal{N}}\times\mathbb{R}^{M\times C\times\mathcal{L}}\to\mathbb{R}^{C}$ as an ensemble model consisting of $M$ trees:
@@
with the trainable parameters $\boldsymbol{w}=(\boldsymbol{w}_{1},\dots,\boldsymbol{w}_{M})$, $\boldsymbol{b}=(\boldsymbol{b}_{1},\dots,\boldsymbol{b}_{M})$, and $\boldsymbol{\pi}=(\boldsymbol{\pi}_{1},\dots,\boldsymbol{\pi}_{M})$ being randomly initialized.
As shown in Equation (4), tree ensembles exhibit permutation invariance when the order of the $M$ trees is rearranged, which is similar to the permutation invariance observed in the hidden neurons of neural networks. However, as discussed in the next section, tree ensembles exhibit several other types of invariance beyond permutation, setting their behavior apart from that of neural networks. In addition to these invariances, there are several key differences between tree ensembles and neural networks. Due to the hierarchical binary tree structure, the influence of each node parameter on the overall model depends on its node position. Moreover, unlike neural networks, tree ensembles lack the concept of activation and intermediate layers. These factors make it challenging to directly apply the matching strategies used for neural networks to achieve LMC.
@@ 3 Invariances Inherent to Tree Ensembles
3.1 Parameter Modification Processes
Figure 3: Invariances inherent to tree ensembles.
First, we clarify what invariances should be considered for tree ensembles. When we consider perfect binary trees, there are three types of invariance:
  • Tree permutation invariance. In Equation (4), the behavior of the function does not change even if the order of the $M$ trees is altered, as shown in Figure 3(a). This corresponds to the permutation of hidden neurons in neural networks, which has been a subject of previous studies on LMC.
  • Subtree flip invariance. When the left and right subtrees are swapped simultaneously with the inversion of the inequality sign at the split, the functional behavior remains unchanged, which we refer to subtree flip invariance. Figure 3(b) presents a schematic diagram of this invariance, which is not found in neural networks but is unique to binary tree-based models. Since $\sigma(-c)=1-\sigma(c)$ for $c\in\mathbb{R}$ due to the symmetry of $\operatorname{sigmoid}$, the inversion of the inequality is achieved by inverting the signs of $\boldsymbol{w}_{m,n}$ and $b_{m,n}$. [24] also focused on the sign of weights, but in a different way from ours. They paid attention to the amount of change from the parameters at the start of fine-tuning, rather than discussing the sign of the parameters.
  • Splitting order invariance. Oblivious trees share parameters at the same depth, which means that the decision boundaries are straight lines without any bends. With this characteristic, even if the splitting rules at different depths are swapped, functional equivalence can be achieved if the positions of leaves are also swapped appropriately as shown in Figure 3(c). This invariance does not exist for non-oblivious perfect binary trees without parameter sharing, as the behavior of the decision boundary varies depending on the splitting order.
Note that MLPs also have an additional invariance beyond just permutation. Particularly in MLPs that employ ReLU as an activation function, the output of each layer changes linearly with a zero crossover. Therefore, it is possible to modify parameters without changing functional behavior by multiplying the weights in one layer by a constant and dividing the weights in the previous layer by the same constant. However, since the soft tree is based on the $\operatorname{sigmoid}$ function, this invariance does not apply. Previous studies [3, 4, 21] have consistently achieved significant reductions in barriers without accounting for this scale invariance. This could be because changes in parameter scale are unlikely due to the nature of optimization via gradient descent. Conversely, when we consider additional invariances inherent to trees, the scale is equivalent to the original parameters.